An electron traveling at relativistic speed moves perpendicular to a magnetic field of 0.20 𝑇. Its path is circular, with a radius of 15 π‘š. Find: (a) the momentum, (b) the speed, and (c) the kinetic energy of the electron. Recall that, in nonrelativistic situations, the magnetic force π‘žπ‘£π΅ furnishes the centripetal force π‘šπ‘£ 2β„π‘Ÿ. Thus, since 𝑝 = π‘šπ‘£ it follows that 𝑝 = π‘žπ΅π‘Ÿ and this relation holds even when relativistic effects are important.Β 


1
Expert's answer
2022-02-13T12:12:43-0500

a)

"qvB=\frac{mv^2}r,"

"qB=\frac{mv}r,"

"p=qBr=4.8\cdot 10^{-19}~\frac{kg\cdot m}s,"

b)

"v=\frac pm=5.3\cdot 10^{11}~\frac ms,"

c)

"E=\frac {mv^2}2=2.5\cdot 10^{-9}~J."